OpenStudy (anonymous):

Prove by induction that for some natural number n: 1 + 3 + 5 +...+ (2n-1) = n^2

OpenStudy (anonymous):

let p(n)=1+2+3+...+(2n-1)=n^2 prove p(1),then assume p(k) is true prove p(k+1) is also true ,then by induction p(n) is also true for all n

OpenStudy (anonymous):

In case a detailed explanation is needed: Establish the base case, for \(n=1\). That is, show that \[1=n^2\] which should be obvious. Now assume the equation holds for \(n=k\), i.e. that \[1+3+\cdots+(2k-1)=k^2\] and use this to show that it holds for \(n=k+1\), i.e. that \[1+3+\cdots+(2k-1)+(2(k+1)-1)=(k+1)^2\] As a hint, the first \(k\) terms of the left side simplify to \(k^2\) by the induction hypothesis. From there, it's a matter of algebraic manipulation.